OpenStudy (anonymous):

solve for x x^2-1=x+1

OpenStudy (anonymous):

You should really try working these out a bit on your own and say where you've gotten stuck. If you just submit all your problems you won't actually be learning anything. \[x^{2} - 1 = x + 1 \rightarrow x^{2}-x = 0 \rightarrow x(x-1) = 0 \rightarrow x \in \left\{ 0,1 \right\}\]

OpenStudy (anonymous):

i am stuck on this one

OpenStudy (anonymous):

Yes, but where did you get stuck?

OpenStudy (anonymous):

im confused on this whole problem can you teach me how to do this?

OpenStudy (anonymous):

Actually my solution is wrong. It should be \[x^2 - 1 = x + 1 \rightarrow x^2 - x -2 = 0 \rightarrow x = (1 \pm \sqrt{9})/2 \rightarrow x \in \left\{ -1,2 \right\} \]

OpenStudy (anonymous):

You move everything to one side of the equal sign and you see that you have a standard quadratic, with a=1, b=-1, c=-2. \[x = (-b \pm \sqrt{b^2 - 4ac})/2a\]